OpenStudy (anonymous):

how do you simplify (3x+10)^2? i mean, what's the next step after that? i know its supposed to be 9x^2+something+100.. can somebody please help? :)

OpenStudy (anonymous):

9x^2+60x+100 :)

OpenStudy (mimi_x3):

(3x+10)(3x+10) 9x^2+60x+100

OpenStudy (anonymous):

wait a min... that is the ans rite?

OpenStudy (mimi_x3):

Yup, it is right (:

OpenStudy (anonymous):

haha thanks mimi!:D

OpenStudy (anonymous):

thanks so much :) whats the rule for that?? :)

OpenStudy (mimi_x3):

Rule!? um idk, I just expand it

OpenStudy (anonymous):

i mean is it (a+b)^2=a^2+2ab+b^2?

OpenStudy (anonymous):

FOIL First Outside Inside Last

OpenStudy (anonymous):

\[x^2 + y^2 = ( x + y) ( x + y )\]\[x^2 - y^2 = ( x + y) ( x - y )\]

OpenStudy (mimi_x3):

I used the FOIL method xD

OpenStudy (anonymous):

yes purpleberry:)

OpenStudy (anonymous):

abdullah ur wrong...

OpenStudy (anonymous):

thanks so much!! :) <3

OpenStudy (mimi_x3):

Abdullah, is right xD

OpenStudy (diyadiya):

(a+b)^2=a^+2ab+b^2 (a-b)^2=a^2-ab+b^2

OpenStudy (anonymous):

x^2 + y^2 does not equal to (x+y)^2

OpenStudy (anonymous):

no prob at all purpleberry:)

OpenStudy (mimi_x3):

Oh yeah, my bad sorry

OpenStudy (anonymous):

im not wrong .. -_-"

OpenStudy (anonymous):

?? that is wrong...

OpenStudy (diyadiya):

you are right but thats not the equation used over here :)

OpenStudy (diyadiya):

@abdullah

OpenStudy (anonymous):

uha .. sorry

OpenStudy (diyadiya):

noproblem:)

OpenStudy (anonymous):

the rule is .. \[(x+y)^2 = x^2 + ( 2xy ) + y^2\]

OpenStudy (diyadiya):

yes

OpenStudy (anonymous):

yup:P

OpenStudy (anonymous):

Diyadiya u r wrong too .. ;p 1-1 lool

OpenStudy (diyadiya):

lol !! i'm not wrong i wrote a and be in place of x and y that doesn't make any difference

OpenStudy (anonymous):

i mean the second one (a-b)^2=a^2-ab+b^2 should be .. (a-b)^2=a^2-2ab+b^2